For each of the following utility functions: Calculate MUx, MUy, MRSx,y; determine whether or not the property of “more is better” is satisfied for both

goods; determine whether or not the marginal utility of x diminishes,
remains constant, or increases as the consumer buys more x; determine whether or not the marginal rate of substitution diminishes, remains constant, or increases as the consumer substitutes x for y along an indifference curve; and sketch the graph of a typical indifference curve.

(a) U(x,y)=(xy)^1/2

I really need help with the indifference curve part of the question!
Answer #1

U = (xy)1/2 = x1/2y1/2

y1/2 = U / (x1/2)

Squaring,

y = (U2) / x [Equation of indifference curve]

Utility remains constant along an indifference curve (IC) while x and y varies along it. So, for purpose of drawing the IC, let us (hypothetically) assume that U = 10. Then,

y = 100 / x

We may develop following data table:

x y
1 100
5 20
10 10
20 5
25 4
50 2
100 1

Graph of IC:

This is the generalized shape of IC for given utility function, which you can draw without any specific U and x-y values.